slope intercept equation m=3, passing through (1, -2)
Answer by swincher4391(1107)

Use point slope: with point (x1,y1) and slope m
y-y1 = m(x-x1)
y-(-2) = 3(x-1)
y+2 = 3x-3
y = 3x-5